# On the Fourier Coefficients of Critical Gaussian Multiplicative Chaos

### v1

@@ 4.2. The oscillatory range
Recall that $r_{n}=\delta\log n$ where $0<\delta<1/4$. By properties (2) and (3) in Section 2, we know that for any pair $\theta_{1},\theta_{2}\in\mathrm{(I)}$, the processes
(4.6) $X_{s}^{(r_{n})}(\theta_{1})\text{ and }X_{s}^{(r_{n})}(\theta_{2})\text{ for }{s\leq t-r_{n}}$
are independent, identically distributed, and independent from the $\sigma$–algebra generated by the field $X_{s}$ up to time $s\leq r_{n}$. Furthermore, the indicator $\mathbf{1}\big\{\theta_{1},\theta_{2}\in\mathcal{G}_{n,t}^{(A)}\big\}$ factors into
$\mathbf{1}\Big\{X_{r_{n}}(\theta_{i})\leq m(r_{n})+A,i\in\{1,2\}\Big\}$
$\times\mathbf{1}\Big\{X_{s}^{(r_{n})}(\theta_{i})\leq U(s)+A,\forall s\in[0,t-r_{n}],i\in\{1,2\}\Big\}$
$:=\mathcal{I}_{\leq r_{n}}\mathcal{I}_{>r_{n}}$
where the $\mathcal{I}_{\leq r_{n}}$ is measurable with respect to (the $\sigma$-algebra generated by) $X_{r_{n}}$, and $\mathcal{I}_{>r_{n}}$ with respect to $(X^{(r_{n})}_{s})_{s\in[0,t-r_{n}]}$.
It follows that for any $(\theta_{1},\theta_{2})$,
(4.7) $\mathbb{E}\Big[e^{\sqrt{2}(X_{t}(\theta_{1})+X_{t}(\theta_{2}))}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in\mathcal{G}_{n,t}^{(A)}\Big\}\Big]$
$=\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))}\mathcal{I}_{\leq r_{n}}\Big]\mathbb{E}\bigg[\prod_{i=1}^{2}e^{\sqrt{2}(X_{t}(\theta_{i})-X_{r_{n}}(\theta_{i}))}\mathcal{I}_{>r_{n}}\bigg]$
and crucially, for $(\theta_{1},\theta_{2})\in(\mathrm{I})$ the rightmost factor is independent of $(\theta_{1},\theta_{2})$ since $\log|\theta_{2}-\theta_{1}|\leq r_{n}-1\leq r_{n}$. By translation invariance of the law of $(X_{s}^{(r_{n})})_{s\geq 0}$, the factor in question equals
$\mathbb{E}\Big[e^{\sqrt{2}X_{t-r_{n}}^{(r_{n})}(0)}\mathcal{I}_{>r_{n}}\Big]^{2}.$
Using $U(s)\leq\sqrt{2}s$,
$\mathcal{I}_{>r_{n}}\leq\mathbf{1}\big\{X_{s}^{(r_{n})}(0)\leq\sqrt{2}s+A,\forall s\in[0,t-r_{n}]\big\},$
and we conclude by Fubini’s theorem that
$\mathbb{E}[\mathcal{C}_{n,t}(\mathrm{I})]$ $\leq e^{-2r_{n}}\iint_{(\mathrm{I})}e^{in(\theta_{2}-\theta_{1})}\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))}\mathcal{I}_{\leq r_{n}}\Big]\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}$
(4.8) $\times{t}e^{-2t+2r_{n}}\mathbb{E}\Big[e^{\sqrt{2}X_{t-r_{n}}^{(r_{n})}(0)}\mathbf{1}\big\{X_{s}^{(r_{n})}(0)\leq\sqrt{2}t+A,\forall s\in[0,t-r_{n}]\big\}\Big]^{2}.$
We begin by showing that the factor in (4.2) is bounded in $t$ and $n$. Under the measure $\mathbb{Q}$ defined through
$\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}={e^{\sqrt{2}X_{t-r_{n}}^{(r_{n})}(0)-\mathbb{E}[X_{t-r_{n}}^{(r_{n})}(0)^{2}]}},$
the mean of $(X_{s}^{(r_{n})}(0),s\in[0,t-r_{n}]$) is shifted by $\sqrt{2}s$ by Girsanov’s theorem. The quantity on the second line in (4.2) is thus bounded by
$\limsup_{t\to\infty}t\cdot\mathbb{P}\Big(X_{s}^{(r_{n})}(0)\leq A,\,\forall s\in[e,t-r_{n}]\Big)^{2}$
$\ll_{A}\limsup_{t\to\infty}\frac{t}{t-r_{n}}=O_{A}(1),$
by Lemma (2.10).
It remains to show that
$\limsup_{t\to\infty}e^{-2r_{n}}\iint_{(\mathrm{I})}e^{in(\theta_{2}-\theta_{1})}\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))}\mathcal{I}_{\leq r_{n}}\Big]\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}=o_{A}\big((\log n)^{-2}\big).$
For fixed $\theta_{1},\theta_{2}$, consider the change of measure given by
(4.9) $\frac{\mathrm{d}\mathbb{Q}_{\Delta}}{\mathrm{d}\mathbb{P}}=e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))-\mathbb{E}[(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))^{2}]}.$
The mean of $X_{r_{n}}(\theta_{1})$ under $\mathbb{Q}_{\Delta}$ is then shifted by
(4.10) $\sqrt{2}\cdot\text{Cov}\Big(X_{r_{n}}(\theta_{1}),X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2})\Big)=\sqrt{2}\big(r_{n}+K(\Delta)\big)$
where $\Delta=\theta_{2}-\theta_{1}$ and $K=K_{r_{n}}$ (cf.  Equation (2.1). The same is true for $X_{r_{n}}(\theta_{2})$ since $K(\Delta)=K(|\Delta|)$. Noting that
$\mathbb{E}[(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))^{2}]=2r_{n}+2K_{r_{n}}(|\theta_{2}-\theta_{1}|),$
the change of variables $(\theta,\Delta)=(\theta_{1},\theta_{2}-\theta_{1})$ thus yields
$e^{-2r_{n}}\iint_{(\mathrm{I})}e^{in(\theta_{2}-\theta_{1})}\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))}\mathcal{I}_{\leq r_{n}}\Big]\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}$
$=2\int_{\Delta_{n}}^{1}e^{in\Delta+2K(\Delta)}\int_{0}^{1}\mathbb{Q}_{\Delta}\mathbf{(}\mathcal{I}_{\leq r_{n}})\mathrm{d}\theta\mathrm{d}\Delta$
$=2\int_{\Delta_{n}}^{1}e^{in\Delta+2K(\Delta)}\mathbb{Q}_{\Delta}\mathbf{(}\mathcal{I}_{\leq r_{n}})\mathrm{d}\Delta$
where $\Delta_{n}=en^{-\delta}$.
We integrate by parts to exploit the cancellation coming from $e^{in\Delta}$. Doing so yields
$\int_{\Delta_{n}}^{1}$ $e^{in\Delta}e^{2K(\Delta)}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\mathrm{d}\Delta$
(4.11) $=\frac{e^{in\Delta}e^{2K(\Delta)}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})}{in}\bigg|^{1}_{\Delta_{n}}-\int_{\Delta_{n}}^{1}\frac{e^{in\Delta}\partial_{\Delta}\big(e^{2K(\Delta)}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\big)}{in}\mathrm{d}\Delta.$
Noting that $\mathbb{Q}_{\Delta}$ is a probability measure, the first term is
$\ll\frac{e^{2K(\Delta_{n})}+e^{2K(1)}}{n}\ll\frac{e^{2r_{n}}}{n}\leq n^{2\delta-1}=o\big((\log n)^{-2}\big).$
By triangle inequality, the second term in (4.11) is
$\ll\frac{1}{n}\sup_{\Delta\in[\Delta_{n},1]}\Big(|K^{\prime}(\Delta)|e^{2K(\Delta)}+e^{2K(\Delta)}\partial_{\Delta}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\Big)$
(4.12) $\ll\frac{e^{2K(\Delta_{n})}}{n}\bigg(\sup_{\Delta\in[\Delta_{n},1]}|K^{\prime}(\Delta)|+\sup_{\Delta\in[\Delta_{n},1]}\partial_{\Delta}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\bigg).$
@@
Using the fact that $e^{2K(\Delta_{n})}\leq n^{2\delta}$ and that $K^{\prime}$ is uniformly bounded on $\Delta_{n},1]$ by
@@
(cf.  Equation (2.1)) we conclude that (4.12) is
(4.14) $\ll n^{3\delta-1}+n^{2\delta-1}\sup_{\Delta\in[\Delta_{n},1]}\partial_{\Delta}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}}).$
Recalling the mean shift in Equation (4.10),
$\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})=\mathbb{P}\Big(X_{i}\leq-\sqrt{2}K(\Delta)-\frac{3}{2\sqrt{2}}\log r_{n}+A,\,\forall i\in\{1,2\}\Big)$
where $X_{1},X_{2}$ are centered Gaussian random variables with mean 0, variance $r_{n}$ and covariance $K(\Delta)$. Equivalently, by Fubini’s theorem, this equals
$\int_{-\infty}^{H(\Delta)}\int_{-\infty}^{H(\Delta)}p_{\Delta}(x_{1},x_{2})\mathrm{d}x_{1}\mathrm{d}x_{2}$
where $p_{\Delta}$ is the density of $(X_{1},X_{2})$ and $H(\Delta)=-\sqrt{2}K(\Delta)-\frac{3}{2\sqrt{2}}\log r_{n}+A$ for simplicity. We differentiate by repeated applications of Leibniz’s rule, whose validity follows from the uniform bounds established later in the proof. This yields
$H^{\prime}(\Delta)\bigg(\int_{-\infty}^{H(\Delta)}p_{\Delta}\big(x_{1},H(\Delta)\big)\mathrm{d}x_{1}+\int_{-\infty}^{H(\Delta)}p_{\Delta}\big(H(\Delta),x_{2}\big)\mathrm{d}x_{2}\bigg)$
(4.15) $+\iint_{(-\infty,H(\Delta)]^{2}}\partial_{\Delta}p_{\Delta}(x_{1},x_{2})\mathrm{d}x_{1}\mathrm{d}x_{2}.$
The boundary terms in (4.2) are equal by symmetry, and are bounded by $|H^{\prime}(\Delta)|=\sqrt{2}|K^{\prime}(\Delta)|$ since
$\int_{-\infty}^{H(\Delta)}p_{\Delta}\big(x_{1},H(\Delta)\big)\mathrm{d}x_{1}=\frac{e^{-H(\Delta)^{2}/2r_{n}}}{\sqrt{2\pi r_{n}}}\mathbb{P}\big(X_{1}\leq H(\Delta)\big|X_{2}=H(\Delta)\big)\leq 1$
uniformly over $\Delta$. As for the double integral in (4.2), differentiating the bivariate Gaussian density and using the triangle inequality gives
$\ll|K^{\prime}(\Delta)|\bigg(\frac{|K(\Delta)|}{(r_{n}^{2}-K(\Delta)^{2})}\mathbb{P}\big(X\leq H(\Delta)\big)+\frac{\mathbb{E}\big[|X_{1}X_{2}|K(\Delta)\mathbf{1}\big\{X\leq H(\Delta)\big\}\big]}{(r_{n}^{2}-K(\Delta)^{2})}$
(4.16) $+\frac{|K(\Delta)|}{(r_{n}^{2}-K(\Delta)^{2})^{2}}\mathbb{E}\big[\mathbf{1}\big\{X\leq H(\Delta)\big\}(K(\Delta)|X_{1}X_{2}|+r_{n}|X_{1}|+r_{n}|X_{2}|)\big]\bigg)$
where $X=\max(X_{1},X_{2})$. We prove that the right-hand side in (4.2) is bounded by a constant times $r_{n}^{2}|K^{\prime}(\Delta)|$ using rather crude bounds. To begin with, we can safely ignore the $(r_{n}^{2}-K(\Delta)^{2})^{-1}$ terms since
$r_{n}^{2}-K(\Delta)^{2}\geq r_{n}^{2}-(r_{n}-1)^{2}\geq 1,\quad\text{ if $|\Delta|>en^{-\delta}=e^{-r_{n}+1}$},\,r_{n}\geq 1.$
Secondly, we bound the probability term in (4.2) as well as the indicators in the expectations therein by one. Combined with the fact that $K(\Delta)\ll r_{n}$ in this range, this yields the simple bound
$\ll r_{n}^{2}|K^{\prime}(\Delta)|\cdot\Big(\mathbb{E}\big[|X_{1}X_{2}|\big]+\mathbb{E}\big[|X_{1}|\big]+\mathbb{E}\big[|X_{2}|\big]\Big).$
The remaining expectations can be estimated using the Cauchy-Schwarz inequality
$\mathbb{E}[|X_{1}X_{2}|]\ll r_{n},\quad\mathbb{E}[|X_{i}|]\ll\sqrt{r_{n}},\,i\in\{1,2\}.$
By combining the estimates above, we may conclude that
$\sup_{\Delta\in[\Delta_{n},1]}\partial_{\Delta}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\ll\sup_{\Delta\in[\Delta_{n},1]}\Big(r_{n}^{3}|K^{\prime}(\Delta)|+|K^{\prime}(\Delta)|\Big)\leq n^{\delta}r_{n}^{3}$
(where we used (4.13) to bound $|K^{\prime}(\Delta)|$). It follows that from (4.14),
(4.17) $\int_{\Delta_{n}}^{1}e^{in\Delta}e^{2K(\Delta)}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\mathrm{d}\Delta\ll r_{n}^{3}n^{3\delta-1}=o\big((\log n)^{-2}\big),$
and in turn that
$e^{-2r_{n}}\iint_{(\mathrm{I})}e^{in(\theta_{2}-\theta_{1})}\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))}\mathcal{I}_{\leq r_{n}}\Big]\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}=o\big((\log n)^{-2}\big).$
4.3. Dominant contribution
@@
Let $A>0$ and $\mathcal{E}_{n,t}^{(A)}$ be as in the statement of Lemma 4.2. Then
$\limsup_{t\to\infty}\mathbb{E}\big[\mathcal{C}_{n,t}\mathrm{(II)}\big]\ll_{A}\frac{1}{(\log n)^{2}}$
for large enough $n$.
@@
(4.18) $\mathcal{F}_{n,t}(\Delta)=te^{-2t}\cdot\mathbb{E}\Big[\mathbf{1}\big\{0,\Delta\in\mathcal{G}_{n,t}^{(A)}\big\}e^{\sqrt{2}(X_{t}(0)+X_{t}(\Delta))}\Big]$

### v2

@@ 4.1. Initial reductions
Let $\Delta=(\theta_{2}-\theta_{1})$. By the change of variables $(\theta,\Delta)=(\theta_{1},\theta_{2}-\theta_{1})$ and Fubini’s theorem,
$\mathbb{E}[\mathcal{C}_{n,t}(\mathrm{I})]$ $\ll te^{-2t}\iint_{(\mathrm{I})}e^{in(\theta_{2}-\theta_{1})}\mathbb{E}\big[e^{\sqrt{2}(X_{t}(\theta_{1})+X_{t}(\theta_{2}))}\mathbf{1}\big\{\theta_{1},\theta_{2},G_{n,t}{(A)}\big\}\big]\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}$
$=te^{-2t}\int_{\Delta_{n}}^{1}e^{in\Delta}\mathbb{E}\big[e^{\sqrt{2}(X_{t}(0)+X_{t}(\Delta))}\mathbf{1}\big\{0,\Delta\in G_{n,t}{(A)}\big\}\big]\mathrm{d}\Delta,$
where in the second line, we used the translation invariance of the law of $X_{t}$ in the following sense
$\big(X_{s}(\theta),X_{s}(\theta+\Delta)\big)_{s\leq t}\overset{d}{=}\big(X_{s}(0),X_{s}(\Delta)\big)_{s\leq t}.$
The next step will be to exploit the cancellation coming from $e^{in\Delta}$ using integration by parts, for which we first need to transform the expectation in the integrand into a tractable function of $\Delta$. We aim to leverage the fact that since $G_{n,t}{(A)}$ only restricts $X_{s}(0)$ and $X_{s}(\Delta)$ at times $s\geq r_{n}$, it should, in principle, only depend on $(X_{s}(0),X_{s}(\Delta))_{s\leq r_{n}}$ through the value of $(X_{r_{n}}(0),X_{r_{n}}(\Delta))$. Furthermore, since $\log|\Delta|^{-1}\leq r_{n}-1$ when $|\Delta|\geq\Delta_{n}$, the integrand’s dependence on $\Delta$ should only figure in the joint density of $(X_{r_{n}}(0),X_{r_{n}}(\Delta))$.
We make this precise under a tilted measure $\mathbb{Q}_{\Delta}$, defined through
(4.6) $\frac{\mathrm{d}\mathbb{Q}_{\Delta}}{\mathrm{d}\mathbb{P}}=e^{\sqrt{2}(X_{r_{n}}(0)+X_{r_{n}}(\Delta))-\mathbb{E}[(X_{r_{n}}(0)+X_{r_{n}}(\Delta))^{2}]}.$
Noting that
$\mathbb{E}[(X_{r_{n}}(0)+X_{r_{n}}(\Delta))^{2}]=2r_{n}+2K_{r_{n}}(|\Delta|),$
(for $K_{r_{n}}$ in (2.1)) we can write
$te^{-2t}\int_{\Delta_{n}}^{1}e^{in\Delta}\mathbb{E}\big[e^{\sqrt{2}(X_{t}(0)+X_{t}(\Delta))}\mathbf{1}\big\{0,\Delta\in G_{n,t}{(A)}\big\}\big]\mathrm{d}\Delta$
$\ll te^{-2t+2r_{n}}\int_{\Delta_{n}}^{1}e^{in\Delta+2K_{t}(\Delta)}\mathbb{E}_{\mathbb{{Q}}_{\Delta}}\big[e^{\sqrt{2}(X_{t-r_{n}}^{(r_{n})}(0)+X_{t-r_{n}}^{(r_{n})}(\Delta))}\mathbf{1}\big\{0,\Delta\in G_{n,t}{(A)}\big\}\big]\mathrm{d}\Delta.$
By properties (2) and (3) in Section 2, we know that for $\Delta\geq\Delta_{n}$, the processes
(4.7) $X_{s}^{(r_{n})}(0)\text{ and }X_{s}^{(r_{n})}(\Delta)\text{ for }{0\leq s\leq t-r_{n}}$
are independent, identically distributed standard Brownian motions. They are also independent from the $\sigma$–algebra generated by the field $X_{s}$ up to time $s\leq r_{n}$, and their distribution is thus unchanged under $\mathbb{Q}_{\Delta}$. By Girsanov’s theorem, the mean of $X_{r_{n}}(0)$ under $\mathbb{Q}_{\Delta}$ is shifted by
(4.8) $\sqrt{2}\cdot\text{Cov}\Big(X_{r_{n}}(0),X_{r_{n}}(0)+X_{r_{n}}(\Delta)\Big)=\sqrt{2}\big(r_{n}+K(\Delta)\big)$
where $K=K_{r_{n}}$, and the same is true for $X_{r_{n}}(\Delta)$ since $K(\Delta)=K(|\Delta|)$.
Using the decomposition
$\mathbf{1}\big\{0,\Delta\in G_{n,t}{(A)}\big\}$ $=\mathbf{1}\big\{X_{r_{n}}(0),X_{r_{n}}(\Delta)\leq U(r_{n})+A\big\}$
$\times\mathbf{1}\big\{X_{s}^{(r_{n})}(0)\leq U(s+r_{n})+A-X_{r_{n}}(0),\forall s\in(0,t-r_{n}]\big\}$
$\times\mathbf{1}\big\{X_{s}^{(r_{n})}(\Delta)\leq U(s+r_{n})+A-X_{r_{n}}(\Delta),\forall s\in(0,t-r_{n}]\big\}$
and the law of total expectation, we therefore have
$\mathbb{E}_{\mathbb{{Q}}_{\Delta}}\Big[e^{\sqrt{2}(X_{t-r_{n}}^{(r_{n})}(0)+X_{t-r_{n}}^{(r_{n})}(\Delta))}\mathbf{1}\big\{0,\Delta\in G_{n,t}{(A)}\big\}\Big]$
$=\mathbb{E}_{\mathbb{{Q}}_{\Delta}}\bigg[\mathbf{1}\big\{X_{r_{n}}(0),X_{r_{n}}(\Delta)\leq U(r_{n})+A\big\}\times$
$\mathbb{E}\Big[e^{\sqrt{2}X_{t-r_{n}}^{(r_{n})}(0)}\mathbf{1}\big\{X_{s}^{(r_{n})}(0)\leq U(s+r_{n})+A-X_{r_{n}}(0),\forall s\in(0,t-r_{n}]\big\}\Big|\,X_{r_{n}}(0)\Big]$
$\mathbb{E}\Big[e^{\sqrt{2}X_{t-r_{n}}^{(r_{n})}(\Delta)}\mathbf{1}\big\{X_{s}^{(r_{n})}(\Delta)\leq U(s+r_{n})+A-X_{r_{n}}(\Delta),\forall s\in(0,t-r_{n}]\big\}\Big|\,X_{r_{n}}(\Delta)\Big]\bigg].$
Performing the outer integral on positions $(X_{r_{n}}(0),X_{r_{n}}(\Delta))$ then shows that the above equals
$f(\Delta):=\iint_{(-\infty,U(r_{n})+A-\sqrt{2}(r_{n}+K(\Delta))]^{2}}p_{\Delta}(u_{1},u_{2})\beta(u_{1})\beta(u_{2})\mathrm{d}u_{1}\mathrm{d}u_{2},$
where
$p_{\Delta}(u_{1},u_{2})=\frac{1}{2\pi\sqrt{r_{n}^{2}-K(\Delta)^{2}}}\exp\!\bigg(\!\!-\frac{1}{2}\frac{r_{n}(u_{1}^{2}+u_{2}^{2})-2K(\Delta)u_{1}u_{2}}{\bigl(r_{n}^{2}-K(\Delta)^{2}\bigr)}\bigg)$
is the density of a pair $(X_{1},X_{2})$ of centered (real) Gaussian random variables satisfying
$\mathbb{E}[X_{1}X_{2}]=K(\Delta),\quad\mathbb{E}[|X_{1}|^{2}]=\mathbb{E}[|X_{2}|^{2}]=r_{n},$
and
$\beta(u):=\mathbb{E}\Big[e^{\sqrt{2}B_{t-r_{n}}}\mathbf{1}\big\{B_{s}\leq U(s+r_{n})+A-u,\forall s\in(0,t-r_{n}]\big\}\Big],$
for $(B_{t})_{t\geq 0}$ a standard Brownian motion. As intended, the dependence on $\Delta$ is now solely contained in $p_{\Delta}$ and in the bounds of integration.
We now turn to our main task, and show that
(4.9) $\limsup_{t\to\infty}te^{-2t+2r_{n}}\int_{\Delta_{n}}^{1}e^{in\Delta+2K(\Delta)}f(\Delta)\mathrm{d}\Delta=o_{A}\big((\log n)^{-2}\big).$
Integration by parts shows that the left-hand side equals
(4.10) $\limsup_{t\to\infty}\,te^{-2t+2r_{n}}\bigg(\frac{e^{in\Delta+2K(\Delta)}f(\Delta)}{in}\bigg|^{1}_{\Delta_{n}}-\int_{\Delta_{n}}^{1}\frac{e^{in\Delta}\partial_{\Delta}\big(e^{2K(\Delta)}f(\Delta)\big)}{in}\mathrm{d}\Delta\bigg).$
We first handle the boundary term by deriving a uniform bound on $f(\Delta)$. Under the change of measure
$\frac{\mathrm{d}\tilde{\mathbb{P}}}{\mathrm{d}\mathbb{P}}=e^{\sqrt{2}B_{t-r_{n}}-(t-r_{n})},$
the mean of $(B_{s})_{s\leq t-r_{n}}$ is shifted by $\sqrt{2}s$ at time $s$ by Girsanov’s theorem, and it follows that
$\beta(u)$ $=e^{t-r_{n}}\tilde{\mathbb{P}}\big(B_{s}\leq U(s+r_{n})+A-u,\forall s\in(0,t-r_{n}]\big)$
$\leq e^{t-r_{n}}\mathbb{P}\big(B_{s}\leq A+\sqrt{2}r_{n}-u,\forall s\in(e,t-r_{n}]\big)$
(4.11) $\ll\big(A+r_{n}+|u|\big)\frac{e^{t-r_{n}}}{\sqrt{t-r_{n}}}$
for any $u\in\mathbb{R}$, by Equation (2.7) and the fact that $U(s)\leq\sqrt{2}s$ for $s>e$. The implicit constant in (4.2) is uniform over $u,t$ and $n$. Using this pointwise bound and the Cauchy-Schwarz inequality, we deduce that
(4.12) $f(\Delta)$ $\leq\frac{e^{2(t-r_{n})}}{t-r_{n}}\mathbb{E}\Big[\big(A+r_{n}+|X_{1}|\big)\big(A+r_{n}+|X_{2}|\big)\Big]\ll_{A}\frac{r_{n}^{2}e^{2(t-r_{n})}}{t-r_{n}}$
for any $\Delta\in[\Delta_{n},1]$, and in turn that
(4.13) $\limsup_{t\to\infty}\,te^{-2t+2r_{n}}\frac{e^{in\Delta+2K(\Delta)}f(\Delta)}{in}\bigg|^{1}_{\Delta_{n}}\ll_{A}\frac{r_{n}^{2}e^{2K(\Delta_{n})}}{n}\ll_{A}\frac{r_{n}^{2}e^{2r_{n}}}{n}.$
Since $r_{n}=\delta\log n$ and $\delta<1/4$, this is clearly $o_{A}((\log n)^{-2})$.
To handle the second term inside the limit in (4.10), we first note that it is
$\ll_{A}te^{-2t+2r_{n}}\frac{1}{n}\Big(\sup_{\Delta\in[\Delta_{n},1]}f(\Delta)|K^{\prime}(\Delta)|e^{2K(\Delta)}+e^{2K(\Delta)}f^{\prime}(\Delta)\Big)$
(4.14) $\ll_{A}\frac{e^{2K(\Delta_{n})}}{n}\bigg(r_{n}^{2}\cdot\sup_{\Delta\in[\Delta_{n},1]}|K^{\prime}(\Delta)|+te^{-2t+2r_{n}}\cdot\sup_{\Delta\in[\Delta_{n},1]}f^{\prime}(\Delta)\bigg)$
@@
for large enough $t$, by the triangle inequality and the uniform bound for $f(\Delta)$ in (4.12). Using the fact that $e^{2K(\Delta_{n})}\leq n^{2\delta}$ and that $K^{\prime}$ is uniformly bounded by
(4.15) $|K^{\prime}(\Delta)|\leq\int_{1}^{e^{r_{n}}}|{k^{\prime}(u\Delta)}|\mathrm{d}u\leq e^{r_{n}}\|k^{\prime}\|_{\infty}\ll n^{\delta},$
on $\Delta_{n},1]$ (cf.  Equation (2.1)), we conclude that (4.2) is
$\ll r_{n}^{2}n^{3\delta-1}+n^{2\delta-1}(te^{-2t+2r_{n}})\sup_{\Delta\in[\Delta_{n},1]}f^{\prime}(\Delta)$
(4.16) $\ll o\big((\log n)^{-2}\big)+n^{2\delta-1}(te^{-2t+2r_{n}})\sup_{\Delta\in[\Delta_{n},1]}f^{\prime}(\Delta).$
It remains to bound $f^{\prime}(\Delta)$ uniformly over $\Delta_{n},1]$. For simplicity, let
$H(\Delta)=U(r_{n})-\sqrt{2}(r_{n}+K(\Delta))+A$
and recall that
$f(\Delta)=\iint_{(-\infty,H(\Delta)]^{2}}p_{\Delta}(u_{1},u_{2})\beta(u_{1})\beta(u_{2})\mathrm{d}u_{1}\mathrm{d}u_{2}.$
Differentiating $f$ by repeated applications of Leibniz’s rule yields
$f^{\prime}(\Delta)$ $=\iint_{(-\infty,H(\Delta)]^{2}}\partial_{\Delta}p_{\Delta}(u_{1},u_{2})\beta(u_{1})\beta(u_{2})\mathrm{d}u_{1}\mathrm{d}u_{2}$
$+H^{\prime}(\Delta)\beta\big(H(\Delta)\big)\bigg(\int_{-\infty}^{H(\Delta)}p_{\Delta}\big(u_{1},H(\Delta)\big)\beta(u_{1})\mathrm{d}u_{1}+$
(4.17) $\hskip 120.45pt\int_{-\infty}^{H(\Delta)}p_{\Delta}\big(H(\Delta),u_{2}\big)\beta(u_{2})\mathrm{d}u_{2}\bigg).$
The boundary terms are equal by symmetry. Using (4.2), we have
$\int_{-\infty}^{H(\Delta)}p_{\Delta}\big(u_{1},H(\Delta)\big)\beta(u_{1})\mathrm{d}u_{1}$
$\ll\frac{e^{t-r_{n}}}{\sqrt{t-r_{n}}}\frac{e^{-H(\Delta)^{2}/2r_{n}}}{\sqrt{2\pi r_{n}}}\mathbb{E}\big[(A+|X_{1}|+r_{n})\big|X_{2}=H(\Delta)\big],$
which, when combined with the following estimates,
(4.18) $|H(\Delta)|\ll_{A}r_{n},$
(4.19) $\mathbb{E}\big[\big(A+|X_{1}|+r_{n}\big)\big|X_{2}=H(\Delta)\big]\leq\big(|H(\Delta)|+r_{n}\big)\ll_{A}r_{n},$
(4.20) $\beta\big(H(\Delta)\big)\ll_{A}\frac{e^{t-r_{n}}}{\sqrt{t-r_{n}}}|H(\Delta)|\ll_{A}\frac{r_{n}e^{t-r_{n}}}{\sqrt{t-r_{n}}},$
@@
allows us to conclude that $f^{\prime}(\Delta)$ is
$\ll_{A}\frac{e^{2(t-r_{n})}}{t-r_{n}}\bigg(r_{n}^{2}n^{\delta}+\iint_{(-\infty,H(\Delta)]^{2}}\partial_{\Delta}p_{\Delta}(u_{1},u_{2})\beta(u_{1})\beta(u_{2})\mathrm{d}u_{1}\mathrm{d}u_{2}\bigg)$
$\ll_{A}\frac{e^{2(t-r_{n})}}{t-r_{n}}\bigg(r_{n}^{2}n^{\delta}+\iint_{\mathbb{R}^{2}}|\partial_{\Delta}p_{\Delta}(u_{1},u_{2})|(A+r_{n}+|u_{1}|)(A+r_{n}+|u_{2}|)\mathrm{d}u_{1}\mathrm{d}u_{2}\bigg).$
Differentiating the bivariate Gaussian density shows that the integral above is bounded by
$\ll\frac{|K^{\prime}(\Delta)\|K(\Delta)|}{(r_{n}^{2}-K(\Delta)^{2})}$
$\times\mathbb{E}\bigg[\Big(A+r_{n}+|X_{1}|\Big)\Big(A+r_{n}+|X_{2}|\Big)\Big(1+\frac{1}{(r_{n}^{2}-K(\Delta)^{2})}+|X_{1}\|X_{2}|\Big)\bigg],$
which is
$\ll|K^{\prime}(\Delta)\|K(\Delta)|\mathbb{E}\Big[\big(A+r_{n}+|X_{1}|\big)\big(A+r_{n}+|X_{2}|\big)\big(1+|X_{1}X_{2}|\big)\Big]$
when $r_{n}\geq 1$, since
$r_{n}^{2}-K(\Delta)^{2}\geq r_{n}^{2}-(r_{n}-1)^{2}\geq 1,\quad\text{ when $|\Delta|>en^{-\delta}=e^{-r_{n}+1}$}.$
Noting that $|K^{\prime}(\Delta)K(\Delta)|\leq r_{n}n^{\delta}$ by Equation (4.15) and that
$\mathbb{E}\Big[\big(A+r_{n}+|X_{1}|\big)\big(A+r_{n}+|X_{2}|\big)\big(1+|X_{1}X_{2}|\big)\Big]\ll_{A}r_{n}^{3}$
by the Cauchy-Schwarz inequality, it follows that
$f^{\prime}(\Delta)\ll_{A}\frac{e^{2(t-r_{n})}}{t-r_{n}}\big(r_{n}^{2}n^{\delta}+r_{n}^{4}n^{\delta}\big),$
and finally that
(4.22) $\limsup_{t\to\infty}n^{2\delta-1}(te^{-2t+2r_{n}})\sup_{\Delta\in[\Delta_{n},1]}f^{\prime}(\Delta)\ll_{A}n^{3\delta-1}r_{n}^{4}=o_{A}\big((\log n)^{-2}\big).$
The bound in (4.9) then follows from Equations (4.13), (4.2), and (4.22).
4.3. Dominant contribution ($|\theta_{2}-\theta_{1}|\leq\Delta_{n}$)
@@
Let $\mathcal{C}_{n,t}(\mathrm{II})$ be as in (4.4). Then for any $\varepsilon>0$,
$\limsup_{t\to\infty}\mathbb{E}\big[\mathcal{C}_{n,t}\mathrm{(II)}\big]=O_{A}\big((\log n)^{-1/2+\varepsilon}\big)$
for large enough $n>0$.
@@
(4.23) $\mathcal{F}_{n,t}(\Delta)=te^{-2t}\cdot\mathbb{E}\Big[\mathbf{1}\big\{0,\Delta\in G_{n,t}{(A)}\big\}e^{\sqrt{2}(X_{t}(0)+X_{t}(\Delta))}\Big]$